Circular No. 7066

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1998ey IN NGC 7080 P. Garnavich, S. Jha, and R. Kirshner, Harvard-Smithsonian Center for

Astrophysics, report: “A spectrum of SN 1998ey was obtained by P. Berlind (Fred L. Whipple Observatory) with the 1.5-m Tillinghast telescope on Dec. 10.1 UT. The spectrum shows four broad peaks at rest wavelengths of 394, 453, 537, and 650 nm (assuming a recession velocity of 4670 km/s, derived from narrow Hα emission from the host galaxy). The widths of features at 537 and 650 nm are 20 000 km/s FWHM. The spectrum is identical to that of the peculiar type-Ic supernova 1997ef (cf. IAUC 6778, 6786), ∼ 2 weeks after discovery, and similar to early spectra of SN 1998bw associated with GRB 980425 (IAUC 6895). A search of the BATSE triggers between Nov. 1 and Dec. 4 shows no clear GRB counterpart to SN 1998ey.”

PKS 0405–385 L. Kedziora-Chudczer, University of Sydney; D. Jauncey, A. Tzioumis,

J. Reynolds, and M. Wieringa, Australia Telescope National Facility; G. Nicholson, Hartebeesthoek Radio Astronomy Observatory; and H. Bignall, University of Adelaide, write: “Four days of observing in our weekly monitoring of the quasar PKS 0405–385 (z = 1.285) with the Australia Telescope Compact Array between Nov. 25 and Dec. 2 shows that rapid, quasiperiodic flux-density variations have reappeared in this unusual source. Such variations were first seen in June 1996 (IAUC 6418), lasting for six weeks. The present variations are smaller in amplitude than in 1996, reaching peak-to-peak values of up to 100, 180, 140, and 60 mJy at wavelengths of 3, 6, 13, and 20 cm. The variations appear also to be slower than in 1996, reaching a maximum of 110 mJy in 3 hr at 6 cm. The inferred brightness temperature is ∼ 3 × 10^{19} K if the variability is intrinsic to the source. PKS 0405–385 shows again the most rapid radio variability yet found.”

COMET C/1998 U5 (LINEAR) K. Dennerl, J. Englhauser, and J. Trümper, Max-Planck-Institut für

Extraterrestrische Physik, Garching; and C. Lisse, University of Maryland, communicate: “Comet C/1998 U5 (cf. IAUC 7059) was observed with the ROSAT Position Sensitive Proportional Counter (PSPC) on Dec. 8– 9. Quicklook data show a clear detection of the comet at soft x-rays (E < 1 keV). For an investigation of the x-ray emission process, observations at other wavelengths are encouraged. Particularly valuable would be estimates of the gas- and dust-production rates.”

1998 December 10  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green